OpenStudy (eric_d):

Expand (1+x)^10 up to and including the term in x^3.Hence,obtain an approximation for (1.01)^10 and (0.99)^10

OpenStudy (eric_d):

Hw shld I begin

ganeshie8 (ganeshie8):

write out the binomial expansion formula

ganeshie8 (ganeshie8):

\[\large (x+1)^{10} = \sum \limits_{k=0}^{10} \binom{10}{k}x^k1^{10-k} \] \[\large = \sum \limits_{k=0}^{10} \binom{10}{k}x^k \]

ganeshie8 (ganeshie8):

expand the sum ^^

OpenStudy (eric_d):

hw to expand

ganeshie8 (ganeshie8):

know how to expand below : \[\large \sum \limits_{k=0}^3 k\] ?

OpenStudy (eric_d):

nope

ganeshie8 (ganeshie8):

its easy : \[\large \sum \limits_{k=0}^3 k \] is just a lazy way of writing : \[\large 0+1+2+3\]

ganeshie8 (ganeshie8):

\[\large \sum \limits_{k=0}^3 k = 0+1+2+3 \]

ganeshie8 (ganeshie8):

based on that, can you guess what below expansion would be : \[\large \sum \limits_{k=0}^3 k^2 = ? \] ?

OpenStudy (eric_d):

0+1+2+3

ganeshie8 (ganeshie8):

think again, you have \(\large k^2\) right ?

OpenStudy (eric_d):

0+1+4+9

OpenStudy (eric_d):

like this

ganeshie8 (ganeshie8):

exactly ! \[\large \sum \limits_{k=0}^3 k^2 = 0^2 + 1^2 + 2^2 + 3^2 \]

ganeshie8 (ganeshie8):

think of left hand side as a one full BOX full of gifts, when u open the BOX, you will see so many gifts(terms on right hand side)

ganeshie8 (ganeshie8):

Alright, lets get back to the original problem : \[\large \sum \limits_{k=0}^{10} \binom{10}{k}x^k\] \[\large = \binom{10}{0}x^0 +\binom{10}{1}x^1 +\binom{10}{2}x^2 +\binom{10}{3}x^3 + \cdots \]

ganeshie8 (ganeshie8):

see if above expansion makes sense.. just find the coefficients and plug in ^^

OpenStudy (eric_d):

yes..

OpenStudy (eric_d):

1

ganeshie8 (ganeshie8):

what/who/when 1 ?

OpenStudy (eric_d):



ganeshie8 (ganeshie8):

find the remaining coefficients also and plug in

OpenStudy (eric_d):